OpenStudy (anonymous):

pure water is poured at the rate of 3 gal/min to a tank containing 300 lbs. of salt dissolved in 100 gals. of water and solution kept while stirring, forced out at 2 gals/min. Find the amount of salt at the end of 1 hr. and 30 minutes?

OpenStudy (dumbcow):

Tank loses 2 gal/min while adding 3 gal/min for a net of 1 gal/min \[Volume = 100 + t\] Salt content in the tank starts at 300 and loses 2C lbs every min where C(t) is the concentration of salt at time t C(t) = Salt content/Volume \[C(t) = \frac{300-2t*C(t)}{100+t}\] solving for C(t) \[\rightarrow C(t) = \frac{300}{100+3t}\] Let S(t) be salt content at time t S(t) = C(t)*Volume \[\rightarrow \ s(t) = \frac{300(100+t)}{100+3t}\] when t=90 \[s(90) = \frac{300(190)}{100+270} = 203.57\]

OpenStudy (anonymous):

thanks for the effort dumb cow, but i already got the answer. . it's 83.10 lbs.

OpenStudy (dumbcow):

Can you please explain to me how you got that answer? Thanks

OpenStudy (anonymous):

this problem is actually related to mixture problem related to differential equation application. . have you encountered this one?

OpenStudy (dumbcow):

no

OpenStudy (anonymous):

here, let me show you how i got it. .

OpenStudy (anonymous):

 we got 100 +t as denominator for the because the rate of going in minus the rate in going out (3t - 2t) resulted to t. .